OpenStudy (anonymous):

Solve for the variable. b/8=3

OpenStudy (anonymous):

b=24

OpenStudy (anonymous):

b/8=3--->b=8*3----b=24

OpenStudy (amistre64):

seeing how giggles is prolly adept at solving this; im glad he provided some sort of steps to aid you :)

OpenStudy (anonymous):

Doing Flvs probably